Suppose Acap Corporation will pay a dividend of $ 2.76 per share at the end of this year and $ 2.94 per share next year. You expect​ Acap's stock price to be $ 53.88 in two years. Assume that​ Acap's equity cost of capital is 11.6 %. a. What price would you be willing to pay for a share of Acap stock​ today, if you planned to hold the stock for two​ years?

b. Suppose instead you plan to hold the stock for one year. For what price would you expect to be able to sell a share of Acap stock in one​ year?

c. Given your answer in ​(b​), what price would you be willing to pay for a share of Acap stock today if you planned to hold the stock for one​ year? How does this compare to your answer in ​(a​)?
Answer #1

A) P0 = $2.76/1.116 + ($2.94+$53.88)/1.1162 = $48.095

B) P1 = (2.94+$53.88)/1.116 = $50.914

C) P0 = ($2.76+$50.914)/1.116 = $48.095

.

PLEASE DO GIVE A THUMBS UP I REALLY NEED IT